At equilibrium, the concentrations in this system were found to be [N2]=[O2]=0.100 M and [NO]=0.500 M. The balanced chemical equation is N2+O2<--->2NO.If more NO is added, bringing its concentration to 0.800 M, what will the final concentration of NO be after equilibrium is re-established?
Answer #1
Concepts and reason

According to Le Chatelier’s principle, in a chemical reaction at equilibrium, if concentration is changed then equilibrium will shift in a direction that oppose the change in concentration. A new equilibrium is established by this change.

Fundamentals

For a reaction,

aA+bB
cC+dD

The equilibrium constant, can be written as

„[a],[v]
„la].[ɔ] – °x

Here,,and are concentration of A, B , C and D respectively.

The balanced chemical equation is,

N2(g)+O2(g)=2NO(8)

Equilibrium constant for this reaction can be written as:

[NO]
[N][0]

Now, substitute 0.500 M for , 0.100 M for and 0.100 M for in expression of equilibrium constant and calculate value of .

1.
(0.500)
(0.100)(0.100)
= 25.0

When more is added then concentration increases to 0.800 M. Then, according to Le Chatelier’s principle, the reaction will shift towards left.

Construct ICE table to calculate change in concentration.

0.800
I(M)
C(M)
E(M)
N2(g) + O2(g) 2NO(g)
0.100 0.100
+x +x
- 2x
0.100+x 0.100+x 0.800 - 2x

Now, equilibrium constant for this reaction is as follow.

Substitute 0.100+x
for and , 0.800 - 2x
for , 25.0 for and calculate value of x.

25 - (0.800 – 2x)
(0.100+x)
5=
0.800-2x
0.100+x
0.500+ 5x =0.800 - 2x
x=0.0429

Now, concentration of from ICE table is as follow.

[NO]=0.800–2x

Substitute 0.0429 for x and calculate final concentration of .

[NO]=0.800-2(0.0429)
=0.714 M

Therefore, after re-establishment of equilibrium final concentration ofis 0.714 M.

Ans:

The final concentration of is 0.714 M.